### v1

# The Physics of Local Optimization in Complex Disordered Systems

Abstract
Limited resources motivate decomposing large-scale problems into smaller, “local” subsystems and stitching together the so-found solutions. We explore the physics underlying this approach and discuss the concept of “local hardness”, i.e., complexity from the local solver perspective, in determining the ground states of both P- and NP-hard spin-glasses and related systems. Depending on the model considered, we observe varying scaling behaviors in how errors associated with local predictions decay as a function of the size of the solved subsystem. These errors stem from global critical threshold instabilities, characterized by gapless, avalanche-like excitations that follow scale-invariant size distributions. Away from criticality, local solvers quickly achieve high accuracy, aligning closely with the results of the more computationally intensive global minimization. These findings shed light on how Nature may operate solely through local actions at her disposal.
A prevalent approach to analyzing large-scale problems involves segmenting them into smaller, “local” components which are then solved, followed by assembling the obtained local optimal solutions in a manner akin to a jigsaw puzzle, thereby yielding meaningful results for the original large-scale problem. This modus operandi is natural when constrained by limited time and computational resources. It is especially common when one assumes — as often happens in problems with an underlying geometric structure (e.g., images of physical systems [1] or proteins [2]) — that these local solutions provide correct and useful information about the larger system. Such local few-body solvers may accurately capture key features of general many-body systems governed by local Hamiltonians, a result known in certain contexts as nearsightedness [3, 4, 5, 6, 7]. Here, we explore physical aspects of this approach for determining global ground states (GSs) of both short- and long-range classical spin-glass systems, a known difficult problem [8, 9] using simple local algorithms.
Similar to other complex systems [10, 11, 12, 13, 14], the challenge posed by spin-glasses results from the combination of disorder and frustration leading to a complex free-energy landscape [15]. According to recent results these systems generically have very low-lying excited states that are radically different from each other and from the GSs  [16, 17, 18]. This phenomenon of system-spanning lowest-energy excitations means that optimal (i.e., GS) subsystem solutions might not necessarily match the optimal full system solution. The occurrence of such states is independent of the classification of the corresponding GS problem in the P or NP classes [19]. Indeed, spin-glasses with GSs obtained with polynomial (P) complexity may also exhibit system-spanning lowest-energy excitations [18]. For instance, this is the case for GSs of two-dimensional (2D) Ising spin-glasses on planar graphs, which can be found in polynomial time [20].
Here, we investigate to which extent local (i.e., subsystem) spin-glass solvers approximate global spin-glass GSs. Understanding such features will help uncover the principles that underlie the effectiveness of local solvers. Our numerical analysis (based on exact GS algorithms) concentrates on Ising spin-glasses with Hamiltonian
$H=-\sum_{\langle ij\rangle}J_{ij}\sigma_{i}\sigma_{j},\;\;\;\sigma_{i}=\pm 1,$ (1)
with couplings $J_{ij}$ drawn from a Gaussian distribution of mean $\overline{J}$ and unit variance. Unless stated otherwise, we set $\overline{J}=0$. Here, the $N$ spins are placed on 2D or 3D lattices with open boundaries and nearest-neighbor interactions (Edwards-Anderson (EA) model [21]) or form a Sherrington-Kirkpatrick (SK) [22] model harboring all-to-all interactions (following a rescaling of $J_{ij}$ by $1/\sqrt{N}$ to guarantee a meaningful thermodynamic limit).
Figure 1: Computational setup for determining the relative spin orientation $\sigma_{i_{0}}\sigma_{j_{0}}$ across the central bond (shown in red). We set the lattice constant to unity. The correct value of $\sigma_{i_{0}}\sigma_{j_{0}}$ is that within the GS of the entire system (comprised of black, light blue, and red bonds). The LSBS computes $\sigma_{i_{0}}\sigma_{j_{0}}$ within the GS of the local subsystem (light blue and red bonds).
Previous related studies attempted to combine multiple local solvers to address global optimization problems [23], fuse local GSs for the construction of global spin-glass instances with planted solutions for benchmarking [24], and develop local spin-glass solvers using gauge-transformation-based deep reinforcement learning [25, 26]. Here, we take on a different perspective and directly investigate to which extent one can predict (parts of) the global GSs based on optimization of local subsystems. To reduce this problem to its bare essentials and minimize boundary effects, we study pairs of neighboring spins (bonds) of the model and introduce a “local single bond solver” (LSBS) that examines these neighboring spins within a local subsystem. We first focus on the GS products $\sigma_{i_{0}}\sigma_{j_{0}}$ for bonds $\langle i_{0}j_{0}\rangle$ which, in the simplest case, are located at the center of the system, cf. Fig. 1.
We are interested to see whether the value of this spin product in the global GS matches that obtained by the LSBS for a subsystem centered on this single bond. Recent studies [27] examined correlations between adjacent spins at the system center after resampling its “shell.” By contrast, here we examine whether this product as determined by (i) GS computations for small subsystems containing this bond and using open boundaries agrees with that in (ii) the global system GS. As we shall see, system-spanning correlated changes to spin-glass GSs appear only at special critical interaction strengths [18]. When the spin-glass couplings are far removed from their critical threshold values, at which global avalanche-like changes occur, the GSs are robust to small variations of the coupling constants. This robustness enables the LSBS to become accurate whenever coupling constants are far from their critical values. We find that for large lattice systems, the LSBS error rate becomes independent of the global system size and exhibits a power-law decay with subsystem size. More generally, several other distance-dependent quantities also show power-law behaviors (see also Ref.[28], LABEL:sec:scaling_subsytem_jc). By contrast, the dependence of the LSBS error rate on the deviations of the coupling from its critical value assumes a stretched exponential form.
Figure 2: The disorder-averaged LSBS error rate $\mathcal{E}_{ij}$ as a function of the subsystem size $L_{\rm sub}$ (resp. $N_{\rm sub}$) for (a) the square-lattice (2D) EA model, (b) the cubic lattice (3D) EA model, and (c) the SK model. For each plot, the results for varying system sizes $L$ ($N$) are shown. For (a), both cases $\overline{J}=0$ and $\overline{J}=1.2$ are shown, whereas in (b) and (c) $\overline{J}=0$. Note the nearly perfect collapse of different system-size data indicating that the error of the LSBS becomes asymptotically independent of system size. For the lattice systems in (a)–(b) and $\overline{J}=0$, the error rate is fit by Eq. (2), cf. the straight dashed lines. Unlike these lattice systems, the all-to-all SK model of (c) is devoid of geometric locality. Here, $N_{\rm sub}$ spins were randomly chosen to comprise the subsystem.
To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [29] based on Blossom V [30] for the polynomial 2D problem and the general-purpose Gurobi software [31] for 3D systems and the SK model. Figure 1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases). If the results agree, the bond product was predicted correctly (with zero error). For $L\times L$ square lattices of sizes $L=256$, $512$, $1024$, we examined $L_{\rm sub}\times L_{\rm sub}$ subsystems with $4\leq L_{\rm sub}\leq 512$, using $\sim 60,000$ disorder realizations. For cubic lattices of side lengths $L=10$ and $12$, the $L_{\rm sub}\times L_{\rm sub}\times L_{\rm sub}$ subsystems were of size $4\leq L_{\rm sub}\leq 10$, and we used $\sim 2,000$ realizations. From these calculations we estimate the LSBS error rate $\mathcal{E}$. Figure 2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases. For the $\overline{J}=0$ lattice systems we find that, although the 2D and 3D spin-glass GS problems belong to different complexity classes (P and NP, respectively), in both cases the error rate is well described by a power-law decay in the subsystem size  [^1] While we will largely focus on the central nearest neighbor bond (with $i=i_{0}$ and $j=j_{0}$), in the End Matter we discuss more general situations in which the sites $i$ and $j$ are a distance ${\sf r}_{ij}>1$ apart. To that end and in order to avoid cumbersome notation, we will often dispense with “0” subscripts.  [^2] In 2D, for the largest values of $L_{\rm sub}={\cal{O}}(L)$, the righthand side of Eq. (2) constitutes an upper bound on the error $\mathcal{E}_{ij}$. Clearly, when $L_{\rm sub}=L$ (when the subsystem is the entire system), the error $\mathcal{E}_{ij}=0$ by definition., viz.
$\mathcal{E}_{ij}\equiv\frac{1-\left[\sigma_{i}\sigma_{j}\sigma^{\rm sub}_{i}\sigma^{\rm sub}_{j}\right]}{2}\sim(\ell_{\mathcal{E}}/L_{\rm sub})^{\kappa}.$ (2)
[…]
$d$ $\ell_{\mathcal{E}}$ $\kappa$ $k_{J}$ $\alpha$
2D 0.20(1) 0.685(8) 10.9(5) 1.35(5)
3D 0.0006(10) 0.18(3) 0.47(3) 1.4(1)
$d$ $\beta$ $\ell_{c}$ $a_{\mathrm{c}}$
2D 0.198(8) 1.01(3) 1.26(1)
3D 1 1.40(9) 1.3(1)
Table 1: Parameters of the fits of the functional forms of Eqs. (2), (3), and (4) to our data. In 2D, $\chi^{2}/\rm{d.o.f}$ is found to be $0.821$, $0.477$, and $1.067$ for Eqs. (2), (3), (4), respectively. In 3D, these values are $1.265$, $0.794$, $0.656$. Note that for the fit of the form (3) in 3D, we fixed $\beta=1$. For the 2D systems, the form (2) was fitted with a cutoff $L_{\rm sub}\geq 16$ in order to minimize finite-size effects.
While the precise power-law decay differs between 2D and 3D, this variability is much less dramatic than maybe expected from the P vs. NP contrast in computational complexity between these cases. This observation suggests a “local hardness” computational complexity descriptor beyond the P-NP classification [19] (and related categories [34]) shedding further light on the intrinsic difficulty of spin-glass GS computations. By local hardness, we allude to how large the subsystem on which computations are done must be so that, when averaged over many instances, the local solver provides a correct answer up to a fixed error rate. Arguably, the local complexity is a relevant descriptor for physical local measurements (the microscopic state of the full many-body system cannot, in general, be probed).
For the case of non-zero average coupling $\overline{J}$, we find the same power-law decay as long as we remain in the (zero-temperature) spin-glass phase, thus suggesting some degree of universality in the behavior of local hardness. As soon as one enters the ferromagnetic phase, however, there is a much faster, exponential decay of the error rate, cf. the data for $\overline{J}=1.2$ in Fig. 2(a), Fig. 7 (End Matter) and further details in Ref. [28], Sec. LABEL:sec:shifted_gaussian. Similarly, we find identical 2D-type scaling and exponents for planar antiferromagnets having faint random perturbations of their couplings from an otherwise constant value (Ref. [28], Sec. LABEL:sec:fully_frustrated). For the all-to-all SK model, on the other hand, the data in Fig. 2(c) reveal that the local hardness does depend on the total system size $N$ and not only on the $N_{\mathrm{sub}}$ randomly chosen spins forming the “local” subsystem, thus indicating a higher degree of local hardness of the SK model. This is consistent with the fact that there is no local geometry in the SK system, and that it consequently features a more complex free-energy landscape with infinitely many thermodynamic states [35]. As a result, one needs to scale $N_{\mathrm{sub}}$ proportional to $N$ to achieve low error rate for the LSBS in the SK model.
We next turn to our central endeavor — that of understanding the physics underlying local solvers. Our approach relies on the concept of “critical thresholds” $J_{\mathrm{c},ij}$ [36, 18], defined as follows. Across any bond, the spin product $\sigma_{i}\sigma_{j}$ is either $+1$ or $-1$. Consider now continuously varying $J_{ij}$ from $\infty$- ∞ to +$\infty$+ ∞ leaving the strengths of all other bonds unchanged. As we have shown elsewhere [18], there is a unique value $J_{ij}=J_{\mathrm{c},ij}$ where the relative orientation of $\sigma_{i}$ and $\sigma_{j}$ changes from $\sigma_{i}=-\sigma_{j}$ ($J_{ij}<J_{\mathrm{c},ij}$) to $\sigma_{i}=\sigma_{j}$ ($J_{ij}>J_{\mathrm{c},ij}$) and, hence, two GSs become degenerate. A similar definition applies for the subsystem — changing the coupling $J_{ij}$ beyond $J_{\mathrm{c},ij}^{\rm sub}$ leads to a change of the subsystem GS. For the LSBS we have the following important observation: If $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ are either (i) both smaller or (ii) both larger than $J_{ij}$, then the LSBS (determining the value of $\sigma_{i}\sigma_{j}$ as calculated within the smaller subsystem GS) will yield the correct $\sigma_{i}\sigma_{j}$ value (i.e., that computed within the global system GS). The opposite case of $J_{ij}$ being sandwiched between $J_{\mathrm{c},ij}$ and $J_{\mathrm{c},ij}^{\rm sub}$, leading to a failure of LSBS, is more likely to occur if $J_{ij}$ is close to the critical thresholds, cf. Fig. 3. Thus, the separation $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ can serve as a predictor of LSBS reliability  [^3] A very similar behavior is observed for the distance $|J_{ij}-J_{\mathrm{c},ij}|$, see the discussions in Supplemental Material Sec. LABEL:sec:scaling_subsytem_jc, but this quantity is not accessible from the perspective of the local solver.. When $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{ij}$ are far apart, then $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ may indeed easily be both larger or both smaller than $J_{ij}$. Conversely, when $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{ij}$ are close, slight miscalculations can lead to $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ to appear on opposite sides of $J_{ij}$, thus causing the LSBS to yield incorrect $\sigma_{i}\sigma_{j}$ values.
[…]
At the critical threshold $J_{c}$, the GS responds sensitively to infinitesimal single-bond changes. Similar to avalanches in the random-field Ising model [38], neural networks [39], systems featuring Highly Optimized Tolerance [40], and other problems [41], the GS spins that are flipped as a result of such changes form fractal boundary Zero Energy Droplets (ZEDs) of sizes conforming to power-law distributions [18]. These distributions may be tied to complex critical spatial correlations (Ref. [28], Sec. LABEL:sec:sg_criticality). Configurations associated with general excitations may be constructed as composites of the (single bond) ZEDs [18]. These excitations are, by construction, of vanishing energy. Right at the thresholds, owing to the sudden sensitivity of the global fundamentally discrete (i.e., Ising) spin-glass GS problem to infinitesimal local changes in the continuous bond couplings, the LSBS achieves low accuracy. However, when the coupling constants are sufficiently removed from these (measure zero) critical thresholds, the local solver becomes progressively more accurate.
Figure 3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS. As is seen here (and in Fig. 4, End Matter), the LSBS error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$. We find this decay to be described well by a stretched-exponential form
$\mathcal{E}_{ij}\sim k_{J}\,e^{-\alpha|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|^{\beta}}.$ (3)
The parameters for $L=1024$, $L_{\rm sub}=8$ square and $L=12$, $L_{\rm sub}=4$ cubic lattice systems appear in Table 1. The modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf. Eq. (LABEL:eq:jccompute) in Ref. [28], Sec. LABEL:sec:compute_jc. Thus the accuracy of the LSBS for a certain bond is determined by the respective excitation energy of the bond in this subsystem.
We further tuned the central bonds of $L=128$ square and $L=12$ cubic lattices to their critical values and then computed the respective critical threshold changes ($\Delta J_{\mathrm{c},ij}$) of all other bonds. We find that the average difference $|\Delta J_{\mathrm{c},ij}|$ decays algebraically with distance from the modified central bond (see Fig. 5, End Matter),
$[|\Delta J_{\mathrm{c},ij}|]\sim(\ell_{c}/r)^{a_{c}}.$ (4)
In 2D and 3D systems, both the length $\ell_{c}$ and the exponent $a_{c}$ are of order unity, the fit parameters are provided in Table 1. For sufficiently large $r$, the central bond coupling negligibly impacts the $J_{\mathrm{c},ij}$ of distant bonds. In other words, the critical coupling of a given bond is almost completely determined only by the couplings on bonds that are close to it. This exemplifies the “locality of the critical threshold $J_{\mathrm{c},ij}$.” Since the ZED volume distribution decays algebraically [18], for sufficiently large distances $r$ between the tuned and observed bond, the probability of ZED excitations connecting these bonds becomes small. When resampling the system boundary, $[|\Delta J_{\mathrm{c},ij}|]$ decays similar to Eq. (4), see Ref. [28], Sec. LABEL:sec:scaling_subsytem_jc. Thus, generating a subsystem by removing peripheral bonds is of limited impact on the critical coupling of the central bond thus explaining the proximity of $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$.
In summary, using exact GS algorithms, we study the effectiveness of local subsystem solvers and relate it to critical threshold ZED physics [18] of system-spanning avalanche-like vanishing-energy excitations. In short-range spin-glasses, away from these special critical thresholds, the response of the system to local perturbations to the GS problem decays rapidly. The difference between local minimization and its full global counterpart becomes significant primarily near the critical coupling threshold values. Away from critical threshold values, local GS solvers achieve increased accuracy. Importantly, the associated error rate of the local solvers becomes independent of the system size [see Fig. 2 (a)] for large systems and decreases rapidly with the subsystem size. Local solvers can thus become increasingly effective for large $L$ where a local solution of a subsystem that is a very small fraction $(L_{\rm sub}/L)^{d}$ of the large system may correctly predict local properties of the large system with negligible error. Simple functions link the average error rate of local solvers to their (subsystem) size as well as to deviations of the coupling constants from their critical threshold value, and changes of the critical coupling of one bond as a result of tuning another distant bond to its critical threshold value. Our analysis reveals that various length-dependent quantities appear to display algebraic scaling just like the scale-free distributions of avalanche size [18], while for ferromagnetic phases we observe an exponential decay of errors with subsystem size (see also End Matter and [28], Sec. LABEL:sec:shifted_gaussian). These behaviors may rationalize the success of more sophisticated local solvers in spin-glasses [24, 25, 26] and other systems [1, 2], and might be broadly applicable in systems not constrained by the “Overlap Gap Property” [42]. For the all-to-all SK model, on the other hand, the error rates always depend on the size of the full system. These differences lead us to conjecture a classification of possible levels of local hardness (Ref. [28], Sec. LABEL:sec:level_hardness). Local solvers are found to be efficient for the considered lattices irrespective of whether the global problems are of P (2D) or NP (3D) complexity. The occurrence of rare events in the form of bonds attaining their critical couplings explains why local solvers cannot in general guarantee to find exact GSs for the full systems.
We briefly speculate on broad implications and extensions of our results regarding the effectiveness of local optimization. Efficient and often highly accurate local minimization is of immense utility in replacing more taxing global (“batch”) gradient descent methods by inherently fast sequential local stochastic gradient descent-type approaches across numerous problems including high-dimensional/large-data machine learning type tasks in many-body physics, materials science, community detection, and other domains [43, 44, 45, 46, 47, 48]. Local plastic variants of the backpropagation algorithm used in machine learning might emulate efficient biological neural computing [49, 50]. Other notable aspects of local minimization and optimized structure in artificial and real neural networks have been further investigated [51, 52]. Locality is of paramount importance in numerous problems in many-body physics [3, 4, 5, 6, 7] where the degrees of freedom may also be continuous. Replacing Eq. (1) by $H=-\sum_{ij}J_{ij}s_{i}s_{j}+u\sum_{i}(s_{i}^{2}-1)^{2}$ with $u\gg 1$ (Eq. (1) corresponds to $u\to\infty$) for continuous $s_{i}\in\mathbb{R}$ suggests that the phenomena that we found might extend to continuous spin systems of sufficiently large quartic interaction strength $u$. In a specific continuous spin variant, that of the spherical model [53, 54], the locality that we described here is altered (Ref. [28], Sec. LABEL:sec:solvable). In future work, we will also further explore viable relations to previously unreported criticality (Ref. [28], Sec. LABEL:sec:sg_criticality). Ultimately, the physics of local optimization may underlie Nature’s enigmatic efficiency, shaping computation across diverse realms.
Acknowledgments. We thank Mark Goh for a discussion in which he alerted us to [42]. Z.N. and M.W. are grateful for support from TU Chemnitz through the Visiting Scholar Program. Z.N. is further grateful for support through the Washington University Seeding Projects for Enabling Excellence & Distinction (SPEED) and the Leverhulme-Peierls Senior Researcher Professorship at Oxford via a Leverhulme Trust International Professorship Grant No. LIP2020-014 (S. L. Sondhi). G.O. gratefully acknowledges support from the Institute for Advanced Study.

### v2

@@
 To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [29] based on Blossom V [30] for the polynomial 2D problem and the general-purpose Gurobi software [31] for 3D systems and the SK model.
-Figure 1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases).
+Fig.
+1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases).
 If the results agree, the bond product was predicted correctly (with zero error).
@@
 From these calculations we estimate the LSBS error rate $\mathcal{E}$.
-Figure 2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases.
+Fig.
+2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases.
 For the $\overline{J}=0$ lattice systems we find that, although the 2D and 3D spin-glass GS problems belong to different complexity classes (P and NP, respectively), in both cases the error rate is well described by a power-law decay in the subsystem size  [^1] While we will largely focus on the central nearest neighbor bond (with $i=i_{0}$ and $j=j_{0}$), in the End Matter we discuss more general situations in which the sites $i$ and $j$ are a distance ${\sf r}_{ij}>1$ apart.
@@
 […]
-$d$ $\ell_{\mathcal{E}}$ $\kappa$ $k_{J}$ $\alpha$
-2D 0.20(1) 0.685(8) 10.9(5) 1.35(5)
-3D 0.0006(10) 0.18(3) 0.47(3) 1.4(1)
+$d$ $\ell_{\mathcal{E}}$ $\kappa$ $k_{J}$ $\sf{J}$
+2D 0.20(1) 0.685(8) 0.198(8) 0.169(7)
+3D 0.0006(10) 0.18(3) 0.47(3) 0.71(5)
 $d$ $\beta$ $\ell_{c}$ $a_{\mathrm{c}}$
-2D 0.198(8) 1.01(3) 1.26(1)
+2D 1.35(5) 1.01(3) 1.26(1)
 3D 1 1.40(9) 1.3(1)
@@
 However, when the coupling constants are sufficiently removed from these (measure zero) critical thresholds, the local solver becomes progressively more accurate.
-Figure 3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS.
-As is seen here (and in Fig.
+Fig.
+3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS.
+As seen here (and in Fig.
 4, End Matter), the LSBS error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$.
-We find this decay to be described well by a stretched-exponential form
-$\mathcal{E}_{ij}\sim k_{J}\,e^{-\alpha|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|^{\beta}}.$ (3)
-The parameters for $L=1024$, $L_{\rm sub}=8$ square and $L=12$, $L_{\rm sub}=4$ cubic lattice systems appear in Table 1.
-The modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf.
+We find this decay to be described well by, respectively, the compressed-exponential/exponential forms
+$\mathcal{E}_{ij}\sim k_{J}\,e^{-(\frac{|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|}{\sf{J}})^{\beta}},$ (3)
+with $\beta>1$ (2D) and $\beta=1$ (3D), see Table 1.
+As in Eq.
+(2), in both 2D and 3D EA systems, $\mathcal{E}_{ij}$ exhibits a similar functional form dependence on the deviation of the coupling of the investigated bond from its critical value (Ref.
+[28], Sec.
+LABEL:sec:2Dvs3D).
+This latter modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf.
 Eq.
@@
 Thus the accuracy of the LSBS for a certain bond is determined by the respective excitation energy of the bond in this subsystem.
-We further tuned the central bonds of $L=128$ square and $L=12$ cubic lattices to their critical values and then computed the respective critical threshold changes ($\Delta J_{\mathrm{c},ij}$) of all other bonds.
+We further tuned the central bonds of the $L=128$ square and $L=12$ cubic lattices to their critical values and then calculated the respective changes in critical threshold ($\Delta J_{\mathrm{c},ij}$) of all other bonds.
 We find that the average difference $|\Delta J_{\mathrm{c},ij}|$ decays algebraically with distance from the modified central bond (see Fig.

### v3

@@
 Limited resources motivate decomposing large-scale problems into smaller, “local” subsystems and stitching together the so-found solutions.
-We explore the physics underlying this approach and discuss the concept of “local hardness”, i.e., complexity from the local solver perspective, in determining the ground states of both P- and NP-hard spin-glasses and related systems.
+We explore the physics underlying this approach and discuss the concept of “local hardness”, i.e., the complexity of predicting local properties of the solution from local information, for the ground-state problem of both P- and NP-hard spin-glasses and related frustrated spin systems.
 Depending on the model considered, we observe varying scaling behaviors in how errors associated with local predictions decay as a function of the size of the solved subsystem.
-These errors stem from global critical threshold instabilities, characterized by gapless, avalanche-like excitations that follow scale-invariant size distributions.
-Away from criticality, local solvers quickly achieve high accuracy, aligning closely with the results of the more computationally intensive global minimization.
-These findings shed light on how Nature may operate solely through local actions at her disposal.
-A prevalent approach to analyzing large-scale problems involves segmenting them into smaller, “local” components which are then solved, followed by assembling the obtained local optimal solutions in a manner akin to a jigsaw puzzle, thereby yielding meaningful results for the original large-scale problem.
+These errors are intimately connected to global critical threshold instabilities, characterized by gapless, avalanche-like excitations that follow scale-invariant size distributions.
+Away from criticality, local solvers quickly achieve high accuracy, aligning closely with the results of the computationally much more expensive global minimization.
+We leverage these findings to introduce a heuristic contraction-based algorithm for globally studying spin-glass ground states.
+The local solvers further display sharp imprints of the phase transition from the spin-glass to the ferromagnetic phase as the distribution of spin-glass couplings is shifted, as well as characteristic differences for the infinite-range model, implying the existence of specific classes of local hardness.
+Our findings shed light on how Nature may operate solely through local actions at her disposal.
+A prevalent, “divide-and-conquer” type, approach to analyzing large-scale problems involves segmenting them into smaller “local” components which are then solved, followed by assembling the obtained local optimal solutions in a manner akin to a jigsaw puzzle, thereby yielding meaningful results for the original large-scale problem.
 This modus operandi is natural when constrained by limited time and computational resources.
-It is especially common when one assumes — as often happens in problems with an underlying geometric structure (e.g., images of physical systems [1] or proteins [2]) — that these local solutions provide correct and useful information about the larger system.
-Such local few-body solvers may accurately capture key features of general many-body systems governed by local Hamiltonians, a result known in certain contexts as nearsightedness [3, 4, 5, 6, 7].
-Here, we explore physical aspects of this approach for determining global ground states (GSs) of both short- and long-range classical spin-glass systems, a known difficult problem [8, 9] using simple local algorithms.
-Similar to other complex systems [10, 11, 12, 13, 14], the challenge posed by spin-glasses results from the combination of disorder and frustration leading to a complex free-energy landscape [15].
-According to recent results these systems generically have very low-lying excited states that are radically different from each other and from the GSs  [16, 17, 18].
-This phenomenon of system-spanning lowest-energy excitations means that optimal (i.e., GS) subsystem solutions might not necessarily match the optimal full system solution.
-The occurrence of such states is independent of the classification of the corresponding GS problem in the P or NP classes [19].
-Indeed, spin-glasses with GSs obtained with polynomial (P) complexity may also exhibit system-spanning lowest-energy excitations [18].
-For instance, this is the case for GSs of two-dimensional (2D) Ising spin-glasses on planar graphs, which can be found in polynomial time [20].
+It is especially common when one assumes — as often happens in problems with an underlying geometric structure (e.g., images of physical systems [Imagenet] or proteins [Alphafold]) — that these local solutions provide correct and useful information about the larger system.
+Such local few-body solvers may accurately capture key features of general many-body systems governed by local Hamiltonians, a result known in certain contexts as nearsightedness [Kohn-near, Cluster1, Cluster2, Cluster3, Cluster4].
+Here, we explore physical aspects of this approach for predicting local as well as global aspects of ground states (GSs) of both short- and long-range classical spin-glass systems, a known difficult problem [cipra2000ising, barahona_computational_1982] using simple local algorithms.
+Similar to other complex systems [SG1, Steinbook, Fortunatobook, Newmanbook, Bryngelson], the challenge posed by spin-glasses results from the combination of disorder and frustration leading to a complex free-energy landscape [janke:07].
+According to recent results these systems generically have very low-lying excited states that are radically different from each other and from the GSs  [Lamarcq, krzakala_disorder_2005, shen2023universal].
+This phenomenon of system-spanning lowest-energy excitations implies that optimal (i.e., GS) subsystem solutions might not necessarily match the optimal full system solution.
+The occurrence of such excitations is independent of the classification of the corresponding GS problem in the P or NP classes [papadimitriou:94].
+Indeed, spin-glasses with GSs of polynomial (P) complexity may also exhibit system-spanning lowest-energy droplets [shen2023universal].
+This is the case for GSs of two-dimensional (2D) Ising spin-glasses on planar graphs, which can be found in polynomial time [bieche:80a].
 Here, we investigate to which extent local (i.e., subsystem) spin-glass solvers approximate global spin-glass GSs.
@@
 Unless stated otherwise, we set $\overline{J}=0$.
-Here, the $N$ spins are placed on 2D or 3D lattices with open boundaries and nearest-neighbor interactions (Edwards-Anderson (EA) model [21]) or form a Sherrington-Kirkpatrick (SK) [22] model harboring all-to-all interactions (following a rescaling of $J_{ij}$ by $1/\sqrt{N}$ to guarantee a meaningful thermodynamic limit).
+Here, the $N$ spins are placed on 2D or 3D lattices with open boundaries and nearest-neighbor interactions (Edwards-Anderson (EA) model [EA]) or form a Sherrington-Kirkpatrick (SK) [SK] model featuring all-to-all interactions (in which case we rescale $J_{ij}$ by $1/\sqrt{N}$ to guarantee a meaningful thermodynamic limit).
 Figure 1: Computational setup for determining the relative spin orientation $\sigma_{i_{0}}\sigma_{j_{0}}$ across the central bond (shown in red).
-We set the lattice constant to unity.
 The correct value of $\sigma_{i_{0}}\sigma_{j_{0}}$ is that within the GS of the entire system (comprised of black, light blue, and red bonds).
 The LSBS computes $\sigma_{i_{0}}\sigma_{j_{0}}$ within the GS of the local subsystem (light blue and red bonds).
-Previous related studies attempted to combine multiple local solvers to address global optimization problems [23], fuse local GSs for the construction of global spin-glass instances with planted solutions for benchmarking [24], and develop local spin-glass solvers using gauge-transformation-based deep reinforcement learning [25, 26].
+Previous related studies attempted to combine multiple local solvers to address global optimization problems [TSM], fuse local GSs for the construction of global spin-glass instances with planted solutions [wang_patch-planting_2017], and develop local spin-glass solvers using gauge-transformation-based deep reinforcement learning [RL_local, RL_local1].
 Here, we take on a different perspective and directly investigate to which extent one can predict (parts of) the global GSs based on optimization of local subsystems.
-To reduce this problem to its bare essentials and minimize boundary effects, we study pairs of neighboring spins (bonds) of the model and introduce a “local single bond solver” (LSBS) that examines these neighboring spins within a local subsystem.
+As a minimal realization of this task we study pairs of neighboring spins (bonds) of the model and introduce a “local single bond solver” (LSBS) that examines neighboring spins within a local subsystem.
 We first focus on the GS products $\sigma_{i_{0}}\sigma_{j_{0}}$ for bonds $\langle i_{0}j_{0}\rangle$ which, in the simplest case, are located at the center of the system, cf.
@@
 We are interested to see whether the value of this spin product in the global GS matches that obtained by the LSBS for a subsystem centered on this single bond.
-Recent studies [27] examined correlations between adjacent spins at the system center after resampling its “shell.” By contrast, here we examine whether this product as determined by (i) GS computations for small subsystems containing this bond and using open boundaries agrees with that in (ii) the global system GS.
-As we shall see, system-spanning correlated changes to spin-glass GSs appear only at special critical interaction strengths [18].
+Recent studies [hartmann_metastate_2023] examined correlations between adjacent spins at the system center after resampling its “shell.” By contrast, here we examine whether this product as determined by (i) GS computations for small subsystems containing this bond and using open boundaries agrees with that in (ii) the global system GS.
+In this way, the LSBS solves a maximally local problem and hence is a diagnostic tool, leading to our definition of “local hardness”.
+As we shall show below, however, these local solutions can be meaningfully stitched together to candidate solutions for the full problems by a contraction-based scheme.
+We will find that system-spanning correlated changes to spin-glass GSs appear only at special critical interaction strengths [shen2023universal].
 When the spin-glass couplings are far removed from their critical threshold values, at which global avalanche-like changes occur, the GSs are robust to small variations of the coupling constants.
@@
 We find that for large lattice systems, the LSBS error rate becomes independent of the global system size and exhibits a power-law decay with subsystem size.
-More generally, several other distance-dependent quantities also show power-law behaviors (see also Ref.[28], LABEL:sec:scaling_subsytem_jc).
-By contrast, the dependence of the LSBS error rate on the deviations of the coupling from its critical value assumes a stretched exponential form.
+Several other distance-dependent quantities also exhibit power-law behaviors (Ref.
+[SM], LABEL:sec:scaling_subsytem_jc).
+By contrast, the dependence of the LSBS error rate on the deviations of the coupling from its critical value is of an exponential form.
 Figure 2: The disorder-averaged LSBS error rate $\mathcal{E}_{ij}$ as a function of the subsystem size $L_{\rm sub}$ (resp.
 $N_{\rm sub}$) for (a) the square-lattice (2D) EA model, (b) the cubic lattice (3D) EA model, and (c) the SK model.
-For each plot, the results for varying system sizes $L$ ($N$) are shown.
+For each plot, the results for varying system sizes $L$ (NN) are shown.
 For (a), both cases $\overline{J}=0$ and $\overline{J}=1.2$ are shown, whereas in (b) and (c) $\overline{J}=0$.
-Note the nearly perfect collapse of different system-size data indicating that the error of the LSBS becomes asymptotically independent of system size.
+Note the nearly perfect collapse of different system-size data, indicating that the error of the LSBS becomes asymptotically independent of system size for (a) and (b), while this is only the case if one scales $N_{\rm sub}$ with NN for the SK model in (c).
 For the lattice systems in (a)–(b) and $\overline{J}=0$, the error rate is fit by Eq.
@@
 the straight dashed lines.
-Unlike these lattice systems, the all-to-all SK model of (c) is devoid of geometric locality.
-Here, $N_{\rm sub}$ spins were randomly chosen to comprise the subsystem.
-To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [29] based on Blossom V [30] for the polynomial 2D problem and the general-purpose Gurobi software [31] for 3D systems and the SK model.
-Fig.
-1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases).
+To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [fractal2d] based on Blossom V [kolmogorov_blossom_2009] for the polynomial 2D problem and the general-purpose Gurobi software [gurobi_optimization_llc_gurobi_2022] for 3D systems and the SK model.
+Figure  1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases).
 If the results agree, the bond product was predicted correctly (with zero error).
-For $L\times L$ square lattices of sizes $L=256$, $512$, $1024$, we examined $L_{\rm sub}\times L_{\rm sub}$ subsystems with $4\leq L_{\rm sub}\leq 512$, using $\sim 60,000$ disorder realizations.
-For cubic lattices of side lengths $L=10$ and $12$, the $L_{\rm sub}\times L_{\rm sub}\times L_{\rm sub}$ subsystems were of size $4\leq L_{\rm sub}\leq 10$, and we used $\sim 2,000$ realizations.
+For $L\times L$ square lattices of sizes $L=256$, $512$, $1024$, we examined $L_{\rm sub}\times L_{\rm sub}$ subsystems with $4\leq L_{\rm sub}\leq 512$, using $\sim 60\,000$ disorder realizations.
+For cubic lattices of side lengths $L=10$ and $12$, the $L_{\rm sub}\times L_{\rm sub}\times L_{\rm sub}$ subsystems were of size $4\leq L_{\rm sub}\leq 10$, and we used $\sim 2\,000$ realizations.
 From these calculations we estimate the LSBS error rate $\mathcal{E}$.
-Fig.
-2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases.
+Figure 2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases.
 For the $\overline{J}=0$ lattice systems we find that, although the 2D and 3D spin-glass GS problems belong to different complexity classes (P and NP, respectively), in both cases the error rate is well described by a power-law decay in the subsystem size  [^1] While we will largely focus on the central nearest neighbor bond (with $i=i_{0}$ and $j=j_{0}$), in the End Matter we discuss more general situations in which the sites $i$ and $j$ are a distance ${\sf r}_{ij}>1$ apart.
 To that end and in order to avoid cumbersome notation, we will often dispense with “0” subscripts.
-[^2] In 2D, for the largest values of $L_{\rm sub}={\cal{O}}(L)$, the righthand side of Eq.
+[^2] In 2D, for the largest values of $L_{\rm sub}={\cal{O}}(L)$, the right-hand side of Eq.
 (2) constitutes an upper bound on the error $\mathcal{E}_{ij}$.
@@
 For the 2D systems, the form (2) was fitted with a cutoff $L_{\rm sub}\geq 16$ in order to minimize finite-size effects.
-While the precise power-law decay differs between 2D and 3D, this variability is much less dramatic than maybe expected from the P vs.
-NP contrast in computational complexity between these cases.
-This observation suggests a “local hardness” computational complexity descriptor beyond the P-NP classification [19] (and related categories [34]) shedding further light on the intrinsic difficulty of spin-glass GS computations.
+While the precise power-law decay differs between 2D and 3D, this variability is much less dramatic than the P vs.
+NP computational complexity contrast between these cases.
+This observation suggests a “local hardness” computational complexity descriptor beyond the P-NP classification [papadimitriou:94] (and related categories [LSA']) shedding further light on the intrinsic difficulty of spin-glass GS computations.
 By local hardness, we allude to how large the subsystem on which computations are done must be so that, when averaged over many instances, the local solver provides a correct answer up to a fixed error rate.
@@
 For the case of non-zero average coupling $\overline{J}$, we find the same power-law decay as long as we remain in the (zero-temperature) spin-glass phase, thus suggesting some degree of universality in the behavior of local hardness.
-As soon as one enters the ferromagnetic phase, however, there is a much faster, exponential decay of the error rate, cf.
+As soon as one enters the disordered ferromagnetic phase, however, there is a much faster, exponential decay of the error rate, cf.
 the data for $\overline{J}=1.2$ in Fig.
@@
 6 (End Matter) and further details in Ref.
-[28], Sec.
+[SM], Sec.
 LABEL:sec:shifted_gaussian.
 Similarly, we find identical 2D-type scaling and exponents for planar antiferromagnets having faint random perturbations of their couplings from an otherwise constant value (Ref.
-[28], Sec.
+[SM], Sec.
 LABEL:sec:fully_frustrated).
@@
 2(c) reveal that the local hardness does depend on the total system size $N$ and not only on the $N_{\mathrm{sub}}$ randomly chosen spins forming the “local” subsystem, thus indicating a higher degree of local hardness of the SK model.
-This is consistent with the fact that there is no local geometry in the SK system, and that it consequently features a more complex free-energy landscape with infinitely many thermodynamic states [35].
-As a result, one needs to scale $N_{\mathrm{sub}}$ proportional to $N$ to achieve low error rate for the LSBS in the SK model.
+This is consistent with the fact that there is no local geometry in the SK system, and that it consequently features a more complex free-energy landscape with infinitely many thermodynamic states [parisi_infinite_1979].
+As a result, one needs to scale $N_{\mathrm{sub}}$ proportionally to $N$ to achieve low error rate for the LSBS in the SK model.
 We next turn to our central endeavor — that of understanding the physics underlying local solvers.
-Our approach relies on the concept of “critical thresholds” $J_{\mathrm{c},ij}$ [36, 18], defined as follows.
-Across any bond, the spin product $\sigma_{i}\sigma_{j}$ is either $+1$ or $-1$.
-Consider now continuously varying $J_{ij}$ from $\infty$- ∞ to +$\infty$+ ∞ leaving the strengths of all other bonds unchanged.
-As we have shown elsewhere [18], there is a unique value $J_{ij}=J_{\mathrm{c},ij}$ where the relative orientation of $\sigma_{i}$ and $\sigma_{j}$ changes from $\sigma_{i}=-\sigma_{j}$ ($J_{ij}<J_{\mathrm{c},ij}$) to $\sigma_{i}=\sigma_{j}$ ($J_{ij}>J_{\mathrm{c},ij}$) and, hence, two GSs become degenerate.
+Our approach relies on the concept of “critical thresholds” $J_{\mathrm{c},ij}$ [newman_ground_2021, shen2023universal], defined as follows.
+Across any bond, the spin product $\sigma_{i}\sigma_{j}$ is either +1+1 or 1-1.
+Consider now continuously varying $J_{ij}$ from -$\infty$ to $\infty$ leaving the strengths of all other bonds unchanged.
+As we have shown elsewhere [shen2023universal], there is a unique value $J_{ij}=J_{\mathrm{c},ij}$ where the relative orientation of $\sigma_{i}$ and $\sigma_{j}$ changes from $\sigma_{i}=-\sigma_{j}$ ($J_{ij}<J_{\mathrm{c},ij}$) to $\sigma_{i}=\sigma_{j}$ ($J_{ij}>J_{\mathrm{c},ij}$) and, hence, two GSs become degenerate.
 A similar definition applies for the subsystem — changing the coupling $J_{ij}$ beyond $J_{\mathrm{c},ij}^{\rm sub}$ leads to a change of the subsystem GS.
 For the LSBS we have the following important observation: If $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ are either (i) both smaller or (ii) both larger than $J_{ij}$, then the LSBS (determining the value of $\sigma_{i}\sigma_{j}$ as calculated within the smaller subsystem GS) will yield the correct $\sigma_{i}\sigma_{j}$ value (i.e., that computed within the global system GS).
-The opposite case of $J_{ij}$ being sandwiched between $J_{\mathrm{c},ij}$ and $J_{\mathrm{c},ij}^{\rm sub}$, leading to a failure of LSBS, is more likely to occur if $J_{ij}$ is close to the critical thresholds, cf.
-Fig.
-3.
-Thus, the separation $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ can serve as a predictor of LSBS reliability  [^3] A very similar behavior is observed for the distance $|J_{ij}-J_{\mathrm{c},ij}|$, see the discussions in Supplemental Material Sec.
+The opposite case of $J_{ij}$ being sandwiched between $J_{\mathrm{c},ij}$ and $J_{\mathrm{c},ij}^{\rm sub}$, leading to a failure of LSBS, is more likely to occur if $J_{ij}$ is close to the critical thresholds.
+This is illustrated in Fig.
+3 correlating the error rate of bond predictions with the couplings’ deviation from criticality.
+Thus, the (subsystem) stability of a bond, captured by the distance $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ between the bare coupling and its subsystem critical threshold can serve as a predictor of LSBS reliability  [^3] A very similar behavior is observed for the distance $|J_{\mathrm{c},ij}-J_{ij}|$, see the discussions in Supplemental Material Sec.
 LABEL:sec:scaling_subsytem_jc, but this quantity is not accessible from the perspective of the local solver..
 When $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{ij}$ are far apart, then $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ may indeed easily be both larger or both smaller than $J_{ij}$.
-Conversely, when $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{ij}$ are close, slight miscalculations can lead to $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ to appear on opposite sides of $J_{ij}$, thus causing the LSBS to yield incorrect $\sigma_{i}\sigma_{j}$ values.
+Conversely, when $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{ij}$ are close, slight miscalculations can lead to $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$ to appear on opposite sides of $J_{ij}$, thus causing the LSBS to yield incorrect $\sigma_{i}\sigma_{j}$ values (see also Ref.
+[SM], Sec.
+LABEL:sec:core_concept).
 […]
-At the critical threshold $J_{c}$, the GS responds sensitively to infinitesimal single-bond changes.
-Similar to avalanches in the random-field Ising model [38], neural networks [39], systems featuring Highly Optimized Tolerance [40], and other problems [41], the GS spins that are flipped as a result of such changes form fractal boundary Zero Energy Droplets (ZEDs) of sizes conforming to power-law distributions [18].
-These distributions may be tied to complex critical spatial correlations (Ref.
-[28], Sec.
+At the critical threshold $J_{\mathrm{c},ij}$, the GS responds sensitively to infinitesimal single-bond changes.
+Similar to avalanches in the random-field Ising model [sethna:01], neural networks [Avalanche1], systems featuring Highly Optimized Tolerance [HOT], and other problems [Avalanche2], the GS spins that are flipped as a result of such changes form what we call Zero Energy Droplets (ZEDs) featuring fractal boundaries and sizes conforming to power-law distributions [shen2023universal].
+In [SM] (Sec.
+LABEL:sec:scaling_kappa) we suggest a simple relation, which appears to be consistent with our data, that links ZED boundary Hausdorff dimensions to the exponent $\kappa$ of Eq.
+(2).
+The ZED distributions and the heterogeneous structure of these rare events, characterized by highly non-uniform spatial distributions, bears resemblance to the temperature [fernandez2013temperature, baity2021temperature, katzgraber_temperature_2007] and disorder chaos [krzakala_disorder_2005, katzgraber_temperature_2007] phenomena in spin glasses, where rare events drive system-spanning changes; these distributions may be tied to complex critical spatial correlations (Ref.
+[SM], Sec.
 LABEL:sec:sg_criticality).
-Configurations associated with general excitations may be constructed as composites of the (single bond) ZEDs [18].
+Configurations associated with general excitations may be constructed as composites of (single bond) ZEDs [shen2023universal].
 These excitations are, by construction, of vanishing energy.
-Right at the thresholds, owing to the sudden sensitivity of the global fundamentally discrete (i.e., Ising) spin-glass GS problem to infinitesimal local changes in the continuous bond couplings, the LSBS achieves low accuracy.
-However, when the coupling constants are sufficiently removed from these (measure zero) critical thresholds, the local solver becomes progressively more accurate.
-Fig.
-3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS.
-As seen here (and in Fig.
-4, End Matter), the LSBS error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$.
-We find this decay to be described well by, respectively, the compressed-exponential/exponential forms
-$\mathcal{E}_{ij}\sim k_{J}\,e^{-(\frac{|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|}{\sf{J}})^{\beta}},$ (3)
-with $\beta>1$ (2D) and $\beta=1$ (3D), see Table 1.
-As in Eq.
-(2), in both 2D and 3D EA systems, $\mathcal{E}_{ij}$ exhibits a similar functional form dependence on the deviation of the coupling of the investigated bond from its critical value (Ref.
-[28], Sec.
+Right at the critical threshold of a given bond, the global GS is extremely sensitive to infinitesimal changes of that local bond coupling, and the LSBS attains its lowest accuracy.
+However, when the coupling constants are sufficiently removed from their (measure zero) critical thresholds, the local solver becomes progressively more accurate.
+As is seen in Fig.
+3 as well as in Fig.
+4, End Matter, the error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$.
+This decay is well described by
+$\mathcal{E}_{ij}\sim k_{J}\,e^{-\left(\frac{|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|}{\sf{J}}\right)^{\beta}},$ (3)
+with $\beta>1$ (2D) and $\beta=1$ (3D), respectively, see Table 1.
+As for the dependence on $L_{\rm sub}$ expressed in Eq.
+(2), the error rate $\mathcal{E}_{ij}$ exhibits the same functional dependence on the ility in 2D and 3D (Ref.
+[SM], Sec.
 LABEL:sec:2Dvs3D).
-This latter modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf.
+The modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the bond excitation energy $\Delta E$ within the subsystem, cf.
 Eq.
 (LABEL:eq:jccompute) in Ref.
-[28], Sec.
+[SM], Sec.
 LABEL:sec:compute_jc.
-Thus the accuracy of the LSBS for a certain bond is determined by the respective excitation energy of the bond in this subsystem.
-We further tuned the central bonds of the $L=128$ square and $L=12$ cubic lattices to their critical values and then calculated the respective changes in critical threshold ($\Delta J_{\mathrm{c},ij}$) of all other bonds.
+Thus, the LSBS accuracy is determined by this excitation energy.
+We further tuned the central bonds of the $L=128$ square and $L=12$ cubic lattices to their critical values and computed the respective changes in critical threshold ($\Delta J_{\mathrm{c},ij}$) of all other bonds.
 We find that the average difference $|\Delta J_{\mathrm{c},ij}|$ decays algebraically with distance from the modified central bond (see Fig.
@@
 In other words, the critical coupling of a given bond is almost completely determined only by the couplings on bonds that are close to it.
-This exemplifies the “locality of the critical threshold $J_{\mathrm{c},ij}$.” Since the ZED volume distribution decays algebraically [18], for sufficiently large distances $r$ between the tuned and observed bond, the probability of ZED excitations connecting these bonds becomes small.
-When resampling the system boundary, $[|\Delta J_{\mathrm{c},ij}|]$ decays similar to Eq.
+This exemplifies the “locality of the critical threshold $J_{\mathrm{c},ij}$.” Since the ZED volume distribution decays algebraically [shen2023universal], for large distances $r$ between the tuned and observed bonds, the probability of ZED excitations connecting these bonds becomes small.
+When resampling the system boundary, $|\Delta J_{\mathrm{c},ij}|]$ decays similar to Eq.
 (4), see Ref.
-[28], Sec.
+[SM], Sec.
 LABEL:sec:scaling_subsytem_jc.
 Thus, generating a subsystem by removing peripheral bonds is of limited impact on the critical coupling of the central bond thus explaining the proximity of $J_{\mathrm{c},ij}^{\rm sub}$ and $J_{\mathrm{c},ij}$.
-In summary, using exact GS algorithms, we study the effectiveness of local subsystem solvers and relate it to critical threshold ZED physics [18] of system-spanning avalanche-like vanishing-energy excitations.
-In short-range spin-glasses, away from these special critical thresholds, the response of the system to local perturbations to the GS problem decays rapidly.
-The difference between local minimization and its full global counterpart becomes significant primarily near the critical coupling threshold values.
-Away from critical threshold values, local GS solvers achieve increased accuracy.
+In summary, using exact GS algorithms, we study the effectiveness of divide-and-conquer type local subsystem solvers and relate it to critical threshold ZED physics [shen2023universal] of system-spanning avalanche-like vanishing-energy excitations.
+In short-range spin-glasses, away from critical thresholds, the response of the system GS to local perturbations decays rapidly.
+Generally, the difference between local and global minimization only becomes significant near critical coupling threshold values.
+Away from these critical values, local GS solvers achieve increased accuracy.
 Importantly, the associated error rate of the local solvers becomes independent of the system size [see Fig.
 2 (a)] for large systems and decreases rapidly with the subsystem size.
-Local solvers can thus become increasingly effective for large $L$ where a local solution of a subsystem that is a very small fraction $(L_{\rm sub}/L)^{d}$ of the large system may correctly predict local properties of the large system with negligible error.
-Simple functions link the average error rate of local solvers to their (subsystem) size as well as to deviations of the coupling constants from their critical threshold value, and changes of the critical coupling of one bond as a result of tuning another distant bond to its critical threshold value.
-Our analysis reveals that various length-dependent quantities appear to display algebraic scaling just like the scale-free distributions of avalanche size [18], while for ferromagnetic phases we observe an exponential decay of errors with subsystem size (see also End Matter and [28], Sec.
+Local solvers may thus become increasingly effective for large $L$.
+Our analysis reveals that various length-dependent quantities display algebraic scaling just like the scale-free distributions of avalanche size [shen2023universal], while for ferromagnetic phases we observe an exponential decay of errors with subsystem size (see also End Matter and [SM], Sec.
 LABEL:sec:shifted_gaussian).
-These behaviors may rationalize the success of more sophisticated local solvers in spin-glasses [24, 25, 26] and other systems [1, 2], and might be broadly applicable in systems not constrained by the “Overlap Gap Property” [42].
+These behaviors may rationalize the success of more sophisticated solvers in spin-glasses [wang_patch-planting_2017, RL_local, RL_local1] and other systems [Imagenet, Alphafold], and might be broadly applicable in systems not constrained by the “Overlap Gap Property” [OGP].
 For the all-to-all SK model, on the other hand, the error rates always depend on the size of the full system.
 These differences lead us to conjecture a classification of possible levels of local hardness (Ref.
-[28], Sec.
+[SM], Sec.
 LABEL:sec:level_hardness).
-Local solvers are found to be efficient for the considered lattices irrespective of whether the global problems are of P (2D) or NP (3D) complexity.
-The occurrence of rare events in the form of bonds attaining their critical couplings explains why local solvers cannot in general guarantee to find exact GSs for the full systems.
-We briefly speculate on broad implications and extensions of our results regarding the effectiveness of local optimization.
-Efficient and often highly accurate local minimization is of immense utility in replacing more taxing global (“batch”) gradient descent methods by inherently fast sequential local stochastic gradient descent-type approaches across numerous problems including high-dimensional/large-data machine learning type tasks in many-body physics, materials science, community detection, and other domains [43, 44, 45, 46, 47, 48].
-Local plastic variants of the backpropagation algorithm used in machine learning might emulate efficient biological neural computing [49, 50].
-Other notable aspects of local minimization and optimized structure in artificial and real neural networks have been further investigated [51, 52].
-Locality is of paramount importance in numerous problems in many-body physics [3, 4, 5, 6, 7] where the degrees of freedom may also be continuous.
-Replacing Eq.
+Local solvers are found to be efficient for the studied lattices irrespective of whether the global problems are of P (2D) or NP (3D) complexity.
+The occurrence of rare events of bonds being close to their critical couplings explains why local solvers cannot, in general, guarantee to find exact GSs of the full systems.
+This mechanism which is reminiscent of chaos in spin glasses [fernandez2013temperature, baity2021temperature, katzgraber_temperature_2007] is generic for the classes of systems we study.
+It is instructive to briefly relate the local-hardness perspective to the metastate and chaotic-size-dependence framework [Newmanbook]: there, one asks whether local observables have an effectively “concentrated” or more “dispersed” distribution over thermodynamic states.
+In our setting, large stability $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ corresponds to a computable proxy for a robust metastate-like concentration of local observables of low LSBS error while small stability reflects competition between nearly degenerate states.
+While these considerations were focused on the hyper-local task of predicting the spins on both ends of a single bond, it is possible to stitch together the results of such calculations to candidate ground states of the full system.
+In the End Matter and Ref.
+[SM] (Sec.
+Contraction solver based on LSBS), we outline an LSBS-based solver that uses the subsystem critical thresholds $J_{\mathrm{c},ij}^{\rm sub}$ to identify “well-determined” bonds, i.e., bonds for which $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is large and the LSBS error rate is exponentially small, cf.
+Eq.
+(3).
+This identification enables the contraction of the system into a smaller number of cohesive large-scale elements (or “communities” [Girvan_Newman2002, Fortunato2010, Comm_detection_local, Comm_detection_local']) before solving for the GS.
+This extension is here presented as proof-of-concept demonstration of how local hardness can guide algorithm design, and further detailed study is required to gauge its full potential.
+Generalized to settings that include external fields and multi-body interactions (Ref.
+[SM], Sec.
+LABEL:sec:multibody), our framework becomes applicable to other important NP-hard optimization problems, such as the Minimum Vertex Cover problem (see End Matter and Ref.
+[SM], Sec.
+LABEL:sec:mvc_ising).
+Locality further underlies numerous many-body physics problems [Kohn-near, Cluster1, Cluster2, Cluster3, Cluster4] where degrees of freedom may also be continuous; replacing Eq.
 (1) by $H=-\sum_{ij}J_{ij}s_{i}s_{j}+u\sum_{i}(s_{i}^{2}-1)^{2}$ with $u\gg 1$ (Eq.
-(1) corresponds to $u\to\infty$) for continuous $s_{i}\in\mathbb{R}$ suggests that the phenomena that we found might extend to continuous spin systems of sufficiently large quartic interaction strength $u$.
-In a specific continuous spin variant, that of the spherical model [53, 54], the locality that we described here is altered (Ref.
-[28], Sec.
+(1) corresponds to $u\to\infty$) for continuous $s_{i}\in\mathbb{R}$ suggests that our findings extend to continuous systems of sufficiently large quartic interaction strength $u$.
+In a continuous spin variant defined by a global constraint– the spherical model [Berlin+Kac, kosterlitz_spherical_1977]– the locality that we described here is altered (Ref.
+[SM], Sec.
 LABEL:sec:solvable).
-In future work, we will also further explore viable relations to previously unreported criticality (Ref.
-[28], Sec.
+Another interesting application area for the ideas outlined here lies in machine learning, with the potential of replacing slow global gradient methods by fast sequential local stochastic gradient descent-type approaches [Harju1997, Comm_detection_local, Comm_detection_local', Ruder2017, review_ML_SGD, MatDis].
+Local plastic variants of the backpropagation algorithm used in machine learning might emulate efficient biological neural computing [Song2020, Goodfellow].
+Other notable aspects of local minimization and optimized structure appear in biological and artificial neural networks [NN_modular, modular_brain].
+We view our LSBS-based contraction solver as a starting point for future investigations of local-hardness–guided heuristics.
+Dividing large scale problems into smaller easier-to-solve fragments may also be beneficial for quantum computing schemes for which smaller scale systems might suffer less from decoherence.
+Extensions of the GS problem analyzed here, including products between distant spins (see Ref.
+[SM], Sec.
+LABEL:sec:distant_spins) to finite temperature correlation functions may provide illuminating connections between our results and finite size scaling [FS1, FS2, FS3, FS4].
+In future work, we will further detail relations to previously unreported criticality (Ref.
+[SM], Sec.
 LABEL:sec:sg_criticality).
-Ultimately, the physics of local optimization may underlie Nature’s enigmatic efficiency, shaping computation across diverse realms.
+The physics of local optimization may underlie Nature’s efficiency, shaping computation across diverse realms.
 Acknowledgments.
-We thank Mark Goh for a discussion in which he alerted us to [42].
+We thank Mark Goh for a discussion in which he alerted us to [OGP].
 Z.N.
@@
 gratefully acknowledges support from the Institute for Advanced Study.
+End Matter
+Here, we provide numerical evidence for Eqs.
+(3) and (4), the existence of a transition in the LSBS accuracy as the distribution of couplings is shifted (so as make the system more ferromagnetic and easier to solve), and we discuss beyond nearest-neighbor correlations found by the LSBS.
+Error rates for deviations from the critical threshold
+Illustrating the applicability of the analysis in the main text, Fig.
+4 demonstrates that the LSBS error indeed decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ in a manner that adheres to the stretched exponential drop of Eq.
+(3).
+Figure 4: Error rate $\mathcal{E}_{ij}$ as a function of $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ for (a) a 2D (square lattice), $L=1024$ system with $L_{\rm sub}=8$, $16$, $32$, $64$ and (b) a 3D (cubic) system of $L=12$ with $L_{\rm sub}=4$, $10$.
+Point percentages denote cumulative probabilities (i.e., fraction of instances) for values of $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ lower than their abscissa.
+Figure 5: Average change of critical threshold values ($|\Delta{J_{\mathrm{c},ij}}|]$) of bonds a distance of $r$ along a Cartesian direction from -$\infty$ to $\infty$ in (a) 2D ($L=128$) and (b) 3D ($L=12$) systems.
+Dashed: Eq.
+(4) with parameters in Table 1.
+As is more evident in 2D, when $r={\cal{O}}(L)$, the average $|\Delta J_{\mathrm{c},ij}|]$ is bounded from above by these ultra-local forms with the fitted $\ell_{c}$ being of the order of the lattice constant.
+.1 Distance dependence of critical threshold values
+As discussed in the main text, the critical threshold of a given bond is largely determined only by the couplings of other bonds that are very close to it.
+When we tuned the central bonds of $L=128$ square and $L=12$ cubic lattices to their critical values, the respective critical threshold changes ($\Delta J_{\mathrm{c},ij}$) of all other bonds at a distance $r$ away conformed to Eq.
+(4) with $\ell_{c}$ of the order of the lattice constant, see Table 1 (with the fit of Eq.
+(4) becoming an upper bound for the largest distances in the 2D system).
+This is illustrated in Fig.
+5.
+In Ref.
+[SM], Sec.
+LABEL:sec:correl_critical_couplings (see Fig.
+LABEL:fig:visualization_Jc_change therein), we further provide a calculated real-space visualization of such a typical rapid drop of the change in the critical coupling away from the origin.
+We further observe a smaller effect on distant bonds for smaller ZEDs.
+Figure 6: For an asymmetric bond distribution, as $\overline{J}$ increases the system transitions from a spin-glass phase to a disordered ferromagnetic phase [melchert_scaling_2009].
+We observe how the fitting coefficient $\kappa$ changes as we adjust $\overline{J}$.
+A larger value of $\kappa$ implies that the error rate of the subsystem solver decays faster, with a crossover to an exponential decay beyond the threshold $1/\overline{J}=0.96$, where the systems enters the ferromagnetic phase [melchert_scaling_2009].
+[…]
+[…]
+Figure 7: Schematic of the contraction solver.
+(a) Well-determined bonds ($|J_{ij}-J_{\mathrm{c},ij}|\geq\vartheta$, where $\vartheta$ is a manually chosen threshold) form connected components that are contracted to effective sites (shown as larger nodes), inheriting all interactions between the component and the rest of the system.
+(b) The relative spin orientations (indicated by ++ and - signs) of sites within each connected component are locked in by the well-determined bonds.
+The GS of the full system can then be reconstructed from the GS of the contracted system by propagating the component orientations.
+Different colors represent distinct connected components.
+Contraction solver based on LSBS
+The LSBS provides local predictions for individual bond spin products.
+Here we outline how these local predictions can be combined into a contraction solver that approximates the full GS spin configuration.
+For each bond, we compute its subsystem critical threshold $J_{\mathrm{c},ij}$ and identify well-determined bonds satisfying $|J_{ij}-J_{\mathrm{c},ij}|\geq\vartheta$, where $\vartheta$ is a chosen threshold.
+These bonds, being far from their critical thresholds, have high local prediction accuracy.
+Well-determined bonds form connected components (or “communities” [Girvan_Newman2002, Fortunato2010, Comm_detection_local, Comm_detection_local']) within which relative spin orientations are reliably fixed.
+Each such component is contracted to a single effective site, inheriting all interactions between the component and the rest of the system.
+The contracted system is solved exactly, and the original spin configuration is reconstructed by propagating the component orientations.
+This contraction procedure is illustrated schematically in Fig.
+7.
+For appropriate choices of $\vartheta$, this approach yields configurations very close to the exact GS while substantially reducing the effective problem size.
+Detailed schematics, algorithmic steps, and performance analysis are provided in the Supplemental Material (Ref.
+[SM], Sec.
+Contraction solver based on LSBS).
+Minimum Vertex Cover as an Ising problem
+As a representative NP-hard optimization problem beyond spin glasses, we also examined the Minimum Vertex Cover (MVC) on random planar graphs generated by Delaunay triangulation.
+Mapping a vertex $i$ being in the cover to $\sigma_{i}=+1$ and otherwise $\sigma_{i}=-1$, the cost function can be written as
+$H_{\text{MVC}}=A\sum_{i\in V}\frac{1+\sigma_{i}}{2}+B\sum_{\langle ij\rangle\in E}\frac{1-\sigma_{i}}{2}\frac{1-\sigma_{j}}{2},$ (5)
+with couplings $J_{ij}=-B/4$ on edges $\langle ij\rangle\in E$ and fields $h_{i}=-A/2+(B/4)\deg(i)$ (for $A=1$, $B=2$).
+Using LSBS on subsystems of topological radius $r$ around a central site, we predict its relative spin orientation in the MVC GS and compare to the exact solution, finding a power-law-like decay of the LSBS error rate with subsystem size (Fig.
+8) with exponent $\kappa\approx 0.8$, close to the spin-glass case.
+This suggests that our locality-based hardness picture applies more broadly to combinatorial optimization problems.
+Figure 8: Minimum Vertex Cover (MVC) as an Ising problem on random planar graphs.
+LSBS error rate for the central spin in the MVC mapping, of the random planar graph shown in the inset, as a function of subsystem size, showing a power-law-like decay with exponent $\kappa\approx 0.8$.
+(Lower left inset) Random planar graph generated from $n=10^{4}$ points by Delaunay triangulation, with sites colored by topological distance $r$ from the geometric center.